an infinetly, long, straight line of charge with pe Cuniform density) E at point P. 2 y WITH COULOMBS
Answer #1

Answer 2 et ma r periai de X Consider a small section q length de on (0,0,2) I dQ=8ed2 2-axis, p.e. de=d2, then differentialE= jde ser da ar 4 17 to (se +2) 3/2 o let 2=r tant limits zot I/2 ..d2=&sec²tdt when 2-0 ty-1/2 and (92+22) 3/2 = {8²+(8 to​​​​​​

Note-please find me in comment box if you find any difficulty. Good luck